OpenStudy (anonymous):

Algebra 2 help! Identify all of the real roots of 2x^3 -42x+40=0

OpenStudy (anonymous):

2x^3-42x+40=0 // divide the left hand side by 2 x^3-42x+40= 0 // according to wolfram alpha the left side of the equation factors to (x-4)(x-1)(x+5) = 0 //therefore x = 4, x= 1, x = -5

OpenStudy (anonymous):

Thank you so much Tomo!